Hello, can someone please help me with this problem? I have no idea how to even go about doing it. Thank you!
"A carpenter charged $110 for a 3 hour job and charged $160 for a 5 hour job. At this rate how much would he charge for a job that required 8 hours to complete?"
Thank you so much for answering my question!

Answer by Alan3354(69443)

"A carpenter charged $110 for a 3 hour job and charged $160 for a 5 hour job. At this rate how much would he charge for a job that required 8 hours to complete?"
-----------
Some assumptions have to be made, but
The difference of 2 hours causes a difference of $50.
That implies a minimum plus an hourly charge.
Hourly = $50/2 = $25/hour
Minimum = $110 - $25*3 = $35
-----------
Charge = $35 + $25*hr
For 8 hours = $35 + $25*8 = $235